A manufacturer wants to increase the shelf life of a line of cake mixes. Past records indicate that the average shelf life of the mix is 216 days. After a revised mix has been developed, a sample of nine boxes of cake mix had a mean of 217.222 days and a standard deviation of 1.2019 days. What is the computed value of the test statistic?
Answer #1

Solution:
Given in the question
Average shelf life of the mix(Population mean) = 216
Sample size (n) = 9
Sample mean = 217.222
Sample standard deviation = 1.2019
Test stat value can be calculated as
Test stat value = (Sample mean - Population mean)/Sample standard deviation/sqrt(n)
Tets stat = (217.222 - 216)/1.2019/sqrt(9) = 1.222/0.4006 = 3.05
So Computed value of test statistic = 3.05